On January 3, 20X1Joe Rockhead, Co. purchased a piece of equipment for $400,000 with a service life of 5 years and a salvage value of $40,000.

Required: Prepare a depreciation schedule for each of the five years under each following assumptions:

(a) straight-line method, and (b) declining balance method at twice the straight-line rate (Double declining balance).
Answer #1

(a) straight-line method,

Year Depreciation expense Accumlated depreciation Book value at the end of year
20X1 72000 72000 328000
20X2 72000 144000 256000
20X3 72000 216000 184000
20X4 72000 288000 112000
20X5 72000 360000 40000
Total 360000

(b) Double decline method,

Straight line rate = 100/5 = 20%

Double decline rate = 20*2=40%

Year Depreciation expense Accumlated depreciation Book value at the end of year
20X1 160000 160000 240000
20X2 96000 256000 144000
20X3 57600 313600 86400
20X4 34560 348160 51840
20X5 11840 360000 40000
Total 360000